I need help please.
The intensity I of light varies inversely w/ the square of the distance d from the source, expressed in the equation I=k/d^2. If the instensity is 1/2 of a candela 16ft from the source, what is the constant of proportionality?
Answer Options:
A. 2 candela-square-feet
B. 32 candela-square-feet
C. 256 candela-square-feet
D. 128 candela-square-feet

Answer by rothauserc(4718)

I=k/d^2. If the instensity is 1/2 of a candela 16ft from the source, what is the constant of proportionality?
solve I = k / d^2 for k by multiplying both sides of = by d^2
k = I * d^2
now I = 1/2 and d = 16, soooooooo
k = 1/2 * 16^2
k = 1/2 * 256
k = 128 candela-square-feet
so the answer is D